OpenStudy (anonymous):

What function best represents a sine function with an amplitude of 4, a period of pi over 6, and a midline at y = −2? f(x) = −2 sin(x − pi over 6) + 4 f(x) = 4 sin 12x − 2 f(x) = 4 sin(x − pi over 6) − 2 f(x) = −2 sin 12x + 4

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

compare the options with below : \[\large f(x) = A\sin(Bx - C) + D\]

ganeshie8 (ganeshie8):

\(|A|\) = Amplitude \(\large \frac{2\pi}{B}\) = Period

ganeshie8 (ganeshie8):

just from the info about amplitude, you can eliminate few options

OpenStudy (anonymous):

Sorry, had to step away from the computer

OpenStudy (anonymous):

A=4 so choice C

ganeshie8 (ganeshie8):

choice C has correct amplitude, but it doesnt have the correct period